# Student reasoning about quantum mechanics while working with physical experiments

### v1

@@ I Introduction
In the work presented here, we investigated how physical experiments may help students improve their conceptual understanding of quantum mechanics by analyzing the in-the-moment student thinking as students worked with the single-photon experiments. We recorded lab sessions of two pairs of students as they worked through a sequence of three experiments that demonstrated both particle-like and wave-like behavior of light. By analyzing the students’ conversations while they worked with the experiments, as well as their reflections on the process afterward, we were able to answer the following research questions:
@@ II.1 Single-photon experiments
The single-photon experiments are currently used in undergraduate courses in a variety of ways to introduce students experimentally to topics such as particle-wave duality of photons and entanglement. They are most often incorporated into either upper-division quantum mechanics or beyond-first-year lab courses. Instructors hope these experiments will help their students accomplish many different learning goals ranging from learning quantum concepts to improving lab skills to increasing motivation for coursework and future research. Based on their goals, instructors choose to incorporate one or several of these experiments, which all use a similar apparatus [15].
@@ II.2 Resources Framework
This study, just like a large portion of physics education research (PER), focuses on students’ conceptual understanding and therefore depends on knowledge of how students learn. One theory of learning is that people learn by actively constructing knowledge themselves [29]. Within this constructionist view, two of the most common approaches taken in PER to determine how instructors can help their students are identifying student misconceptions and identifying the pieces of knowledge or “resources” students use to construct knowledge [30, 31]. Many PER studies focus on discovering student difficulties, which can help instructors know where educational interventions may be most beneficial [30]. Much of that work focuses on finding student misconceptions, so instructors can help students elicit these incorrect ideas, confront evidence demonstrating how the ideas are inaccurate, and then resolve any contradictions in their reasoning [32]. Another approach is to focus on the resources students have, so instructors can help students build off of their own ideas to learn new topics [33, 34]. Instructors’ role is then to help students identify potentially useful ideas they already have and know how and when to apply them [35].
There has been various terminology in the literature used to describe student resources or pieces of knowledge students can use to construct new knowledge (e.g., Refs. [36, 37, 34]), each with its own definition. In this work, we use the term resources, which is popular in the PER community and can encompass any size idea held by an individual student that can be used as a building block to construct additional knowledge [34, 31]. Resources are ideas that students may have learned at any point in their lives, either inside or outside of the classroom, and can be organized in such a way that they even have sub-structure [38]. They may be activated in different contexts, and do not necessarily need to be correct as long as they can be productive in at least one context, leading to the possibility of students holding seemingly contradictory ideas [39]. Once resources have been identified, helping students learn involves helping them to refine the way different resources are activated and organized so that their ideas line up with the canonical understanding of physics [40, 41]. Using the resources framework to study student learning allows us to focus our attention on the students, place value on the knowledge that they have, and understand the diverse ways different students may engage with a single context [31].
The resources and misconceptions frameworks both have characteristics that have been identified in student reasoning, so it can be useful for instructors to consider multiple theoretical models [35]. Especially for quantum mechanics, which is viewed by students as being particularly unintuitive [42], it may be helpful to understand not just what misconceptions students may have, but also what resources they can learn to activate to reason through these complex and unintuitive topics.
@@ II.3 Conceptual understanding of quantum mechanics through an experimental context
Discussing the context of single-photon experiments has helped instructors and researchers identify specific student difficulties and elucidate student ideas surrounding the behavior and properties of photons. Students have been found to have difficulty reasoning about single-photon interference with a Mach-Zehnder interferometer, in particular because they often ignore both that single photons interfere with themselves and the wave-like properties of photons [54, 55]. Other studies have investigated more nuanced student reasoning about the ontology of photons, finding that students classify them in a variety of ways involving particle-like descriptions, wave-like descriptions, combinations of the two, or neither [56, 57, 58, 59, 60, 61]. These student ideas can change between contexts [50, 62] and are affected by instruction [48], including both the words [62] and visualizations [5] used by instructors.
@@
Various kinds of classroom activities have been shown to help students improve their conceptual understanding related to the ideas of particle-wave duality of photons and single-photon interference. One of the simplest ways to have students engage with these ideas in a concrete context is to discuss what would happen in an experiment without the students interacting with an actual experimental apparatus. Classroom discussions of the single-photon experiments as thought experiments have been shown to help students distinguish between the ways uncertainty manifests in classical versus quantum models [52]. In order to provide students opportunities to see how experimental results depend on various parameters, interactive simulations, including of a Mach-Zehnder interferometer with single photons, have been developed and incorporated into some courses. These have been shown to improve students’ conceptual understanding and reduce some of the known student difficulties [6, 8, 55, 9]. Videos of real data combined with diagrams of and questions about the experiment have also been shown to help students engage with quantum reasoning about probability and interference [63], and interactive screen experiments have led to the improvement of students’ understanding of the properties and behavior of photons and the probabilistic interpretation of quantum mechanics [11].
@@
To date, there have been only a few studies investigating how working with the physical apparatus of the single-photon experiments can improve student conceptual understanding of quantum mechanics. Some of the instructors who have developed and published about the use of these experiments in their courses have shown that the experiments help students be motivated to learn about the topics [22], understand the concept of quantum superpositions [19], and correctly answer conceptual questions about entanglement and single-photon interference [14]. Our earlier work that investigates student learning outcomes in courses across many different institutions also showed that students self-report learning concepts in addition to discussing how these experiments helped confirm their belief that quantum mechanics describes the physical world [20]. However, all of the education research performed on students working with the physical experiments has focused on student outcomes instead of the process, thereby missing out on understanding what specific activities prompt productive student reasoning that can lead to various learning gains.
@@ III.1 The experiments
The lab guides the students worked through in our study were designed to align with the way many instructors implement the single-photon experiments in their quantum or beyond-first-year lab courses [15]. We developed the lab guides by looking at examples in the literature and from instructors who had shared their materials with us. We then adjusted the materials to match our logistical constraints, considering the equipment we had available, the amount of time we could ask students to commit, and the way we were implementing these experiments outside of a course setting. The lab guides included metacognitive scaffolding asking the students to reflect with their lab partner on the experimental results. Prior to providing the lab guides to the students, we tested each of the lab guides twice with colleagues who had a Master’s or PhD in a field of physics or astronomy outside of Atomic, Molecular, and Optical (AMO) physics.
All three of the experiments performed by the students utilized a similar apparatus, as shown schematically in Fig. 1. Each began with a 405 nm laser illuminating a non-linear crystal. Inside the crystal, a small fraction of the photons from the laser were converted into pairs of spatially-entangled lower-energy photons through the process of SPDC. These photons were then detected at either two or three detectors, labeled A, B, and B’. The photons arriving at these detectors were converted into electronic signals, which were sent through a set of electronics that generated the number of coincidence counts for all combinations of detectors. Coincidence counts are the number of times photons arrived at each of the detectors within the same very short time window. The single and coincidence counts were displayed on a LabVIEW computer interface. The lab guides provided to the students are detailed in Ref. [64] and summarized below.
@@
The second experiment was a continuation of the first. When the students arrived, the laser was already aligned through the crystal, and detectors A and B’ were already aligned with the down-converted photons. The students began by placing a beam splitter in front of detector B’ to split the path of the photons going to that detector, and then added a fiber coupler for the third detector (detector B) to detect the photons that were reflected at the inserted beam splitter. The students then measured the two- and three-way coincidence counts and used those to calculate the second-order correlation parameter, , of the two outputs of the beam splitter. This allowed them to determine whether the light exiting the two outputs of the beam splitter was correlated (as it would be for a classical model of light, in which the amplitude of the wave splits) or anti-correlated (as it would be for single photons that can be detected at only one of the two outputs). The students obtained a value for the three-detector correlation parameter that fell within the quantum regime. This indicates that when they used detector A to know that they had a single-photon state, they could measure it at only one of the two outputs of the beam splitter. However, when they ignored the counts at detector A, their value fell within the classical regime, since they were not enforcing the fact that they had a single-photon state.
@@
In between the second and third experiments, V.B. added additional optical elements to create and align a Mach-Zehnder interferometer in one of the paths of the down-converted photons. The students began the third experiment by looking at interference with a visible laser (referred to as the “alignment laser”) that was already aligned along that same path. The students already had some familiarity with the alignment laser since they had also used it to help align the beam splitter in the second experiment. The students then spent the rest of the experiment working with the down-converted photons, with which they were immediately able to see interference. To better understand what they were seeing, the students had opportunities to play around with this setup by blocking and unblocking the two arms of the interferometer and ramping the piezoelectric actuator (piezo) that was attached to one of the mirrors in the interferometer. By using a computer-aided control system to apply different voltages to the piezo, the students were able to change the relative path length between the two arms of the interferometer by a distance on the nanometer scale, thereby causing the interference pattern to shift. To finish off the third experiment, the students placed halfwave plates in both arms of the interferometer and rotated one so that the polarizations of the photons in the two arms were orthogonal, thus removing the interference. When they then placed a polarizer (aligned at to the light in both arms of the interferometer) after the second beam splitter, the interference pattern reappeared. This is called a quantum eraser since the which-path information has been “erased.”
@@ III.2 Data sources
Our primary data source is think-aloud lab sessions where each pair of students worked together through a sequence of three two-hour lab sessions while being prompted to discuss their thinking out loud. During these sessions, V.B. acted as both a researcher (explaining the purpose and process of the research and prompting the students to explain their reasoning if needed) and an instructor or teaching assistant (answering questions the students had and ensuring they followed safety protocols). Since she had set-up the experimental apparatus and designed the lab guides, she was familiar with the experiments in a similar way to a typical instructor. All of the lab sessions were video and audio recorded.
After the students had completed all three experiments, they also participated in semi-structured individual interviews over Zoom, which we refer to as “post-interviews.” These lasted approximately one hour and included questions about the concepts covered in these experiments; where in the sequence of experiments the students had learned about these concepts; and other potential learning outcomes such as interest, self-efficacy, and feeling that they had seen real quantum effects [20]. We asked demographic questions at the end of the post-interviews and found that three of the students identified as white and the fourth as white and Asian 111The student specified a specific country in Asia when asked about their race, but to help preserve their anonymity, we have chosen to present only a more general racial categorization here. Two of the students identified as women and two as men, although we use gender neutral pseudonyms and pronouns throughout the paper to help protect their anonymity. The students were compensated with a gift card for the time they spent in the lab sessions and interview.
@@ III.3 Data analysis
Since the students spent the most time sense-making in the third experiment, we focused our analysis of resources solely on that experiment and therefore students’ understanding of single-photon interference. We began to identify student resources by watching the video recordings of the relevant conversations we had noted in the content logs. To avoid missing additional moments, we fully transcribed the third lab session for both groups. After carefully going through the transcripts and returning to watch the video clips of any moments where it seemed information may be missing (e.g., when students were using hand gestures), key conversations were chosen that demonstrated evidence of the identified resources. These were matched with student quotes from the post-interviews where the students reflected on the knowledge they had relied on while working with the experiments.
@@ III.4 Limitations
Although we modeled our lab sessions after those used in actual courses, we investigated only one specific implementation of the single-photon experiments and it was in a non-class setting. The materials we designed were based off of materials used in other courses, but still had to be altered to fit within our specific constraints. Additionally, the context of a research setting is different than a classroom since V.B. was with the students at all times (whereas teaching assistants or instructors often rotate through groups), the students did not receive additional instruction beyond the lab sessions (for example, detailed derivations of the math demonstrated in the experiment were only available if students decided to look up the references in their own time), and the students did not need to consider grades or lab write-ups.
@@ IV Results: Student resources for understanding single-photon interference
Throughout the third lab session (the single-photon interferometer experiment), the students were compelled to make sense of the experimental data that simultaneously demonstrated particle-like and wave-like properties of single photons. This sense-making process consisted of the students discussing their observations with their lab partners and collectively building on various resources the students came in with until they had generated an explanation with which they were satisfied. Using the transcripts of the third think-aloud lab sessions, as well as the post-interviews, we identified four resources the students utilized:
Knowledge of classical wave interference
Knowledge of wave functions
Probabilities represent possible quantum outcomes
Thinking in terms of information.
By Thinking in terms of information, we are referring to the way information (in this case, which-path information) can be used as a quantifiable object to understand a physical system.

### v2

@@ I Introduction
In the work presented here, we investigated how physical experiments may help students improve their conceptual understanding of quantum mechanics by analyzing the in-the-moment student thinking as students worked with the single-photon experiments. We recorded lab sessions of two pairs of students outside of a course context as they worked through a sequence of three experiments that demonstrated both particle-like and wave-like behavior of light. By analyzing the students’ conversations while they worked with the experiments, as well as their reflections on the process afterward, we were able to answer the following research questions:
@@ II.1 Single-photon experiments
The single-photon experiments are currently used in undergraduate courses in a variety of ways to introduce students experimentally to topics such as particle-wave duality of photons and entanglement. They are most often incorporated into either upper-division quantum mechanics or beyond-first-year (BFY) lab courses. When used in BFY lab courses, the students may or may not have already taken a quantum mechanics course. Instructors hope these experiments will help their students accomplish many different learning goals ranging from learning quantum concepts to improving lab skills to increasing motivation for coursework and future research. Based on their goals, instructors choose to incorporate one or several of these experiments, which all use a similar apparatus [15].
@@ II.2 Resources Framework
This study, just like a large portion of physics education research (PER), focuses on students’ conceptual understanding and therefore depends on knowledge of how students learn. One theory of learning is that people learn by actively constructing knowledge themselves [29]. Within this constructionist view, two of the most common approaches taken in PER are identifying student difficulties [30] and identifying the pieces of knowledge or “resources” students use to construct knowledge [31]. The studies focused on discovering student difficulties can help instructors know where educational interventions may be most beneficial and aid in the development of new materials [30, 32, 33, 34]. Another approach is to focus on the resources students have, so instructors can help students build off of their own ideas to learn new topics [35, 36]. Instructors’ can then help students identify potentially useful ideas they already have and learn how and when to apply them [37].
There has been various terminology in the literature used to describe student resources or pieces of knowledge students can use to construct new knowledge (e.g., Refs. [38, 39, 36]), each with its own definition. In this work, we use the term resources, which is popular in the PER community and can encompass any size idea held by an individual student that can be used as a building block to construct additional knowledge [36, 31]. Resources can be broad and encompass many other resources (e.g., “coordinate systems”) [40] or be more specific, around the size of a typical course learning goal (e.g, “forces influence the motion of objects”) [41]. Resources are ideas that students may have learned at any point in their lives, either inside or outside of the classroom [40]. They may be activated in different contexts, and do not necessarily need to be correct as long as they can be productive in at least one context, leading to the possibility of students holding seemingly contradictory ideas [42]. Once resources have been identified, instructors can help students refine the way different resources are activated and organized so that their ideas line up with the canonical understanding of physics [41, 43]. Using the resources framework to study student learning allows us to focus our attention on the students, place value on the knowledge that they have, and understand the diverse ways different students may engage with a single context [31].
The resources and difficulties frameworks both have characteristics that have been identified in student reasoning, so it can be useful for instructors to consider multiple theoretical models [37]. Most of the work investigating student reasoning in the context of the single-photon experiments has focused on student difficulties (e.g., Refs. [44, 45]), so it is also important to identify resources students activate as they reason through these complex and unintuitive topics.
@@ II.3 Conceptual understanding of quantum mechanics through an experimental context
Discussing the context of single-photon experiments has helped instructors and researchers identify specific student difficulties and elucidate student ideas surrounding the behavior and properties of photons. Students have been found to have difficulty reasoning about single-photon interference with a Mach-Zehnder interferometer [44, 45]. In particular, students often ignore the wave-like properties of photons, instead discussing them as point particles traversing the arms of the interferometer. Students also often do not account for interference or the relative phase shift between the two arms of the interferometer when discussing the number of photons that will be detected after the interferometer [44, 45]. Other studies have investigated nuanced student reasoning about the ontology of photons, finding that students classify photons in a variety of ways, including particle-like descriptions, wave-like descriptions, combinations of the two, or neither [57, 58, 59, 60, 61, 62]. These student ideas can change between contexts [53, 63] and are affected by instruction [51], including both the words [63] and visualizations [5] used by instructors.
@@
Various kinds of classroom activities have been shown to help students improve their conceptual understanding related to the ideas of particle-wave duality of photons and single-photon interference. One of the simplest ways to have students engage with these ideas in a concrete context is to discuss what would happen in an experiment without the students interacting with an actual experimental apparatus. Classroom discussions of the single-photon experiments as thought experiments have been shown to help students distinguish between the ways uncertainty manifests in classical versus quantum models [55]. In order to provide students opportunities to see how experimental results depend on various parameters, interactive simulations, including of a Mach-Zehnder interferometer with single photons, have been developed and incorporated into some courses. These have been shown to improve students’ conceptual understanding of single photon interference [6, 8] and reduce the known student difficulty of ignoring the interference of single photons [45, 9]. Videos of real data combined with diagrams of and questions about the experiment have also been shown to help students use the concept of superpositions to explain single-photon interference [64], and interactive screen experiments have led to the improvement of students’ understanding of the properties and behavior of photons and the probabilistic interpretation of quantum mechanics [11].
@@
To date, there have been only a few studies investigating how working with the physical apparatus of the single-photon experiments can improve student conceptual understanding of quantum mechanics. Some of the instructors who have developed and published about the use of these experiments in their courses have shown that the experiments help students be motivated to learn about the topics [22], self-report an improved understanding of quantum superpositions [19], and correctly answer conceptual questions about entanglement and single-photon interference, such as if interference can be observed when photons in the two arms have different polarizations [14]. Our earlier work that investigates student learning outcomes in courses across many different institutions also showed that students self-report learning quantum concepts while working with these experiments [20]. Additionally, the experience allowed many students to confirm their belief that quantum mechanics describes the physical world even as different students provided varied responses as to what is quantum about the experiments [20]. However, all of the education research performed on students working with the physical experiments has focused on student outcomes instead of the process, thereby missing out on understanding what specific activities prompt productive student reasoning that can lead to various learning gains.
@@ III.1 The experiments
The lab guides the students worked through in our study were designed to align with the way many instructors implement the single-photon experiments in their quantum or BFY lab courses [15]. We developed the lab guides by looking at examples in the literature and from instructors who had shared their materials with us. We then adjusted the materials to match our logistical constraints, considering the equipment we had available, the amount of time we could ask students to commit, and the way we were implementing these experiments outside of a course setting. The lab guides included metacognitive scaffolding asking the students to reflect with their lab partner on the experimental results. Prior to providing the lab guides to the students, we tested each of the lab guides twice with colleagues who had a Master’s or PhD in a field of physics or astronomy outside of Atomic, Molecular, and Optical (AMO) physics.
All three of the experiments performed by the students utilized a similar apparatus, as shown schematically in Fig. 1. Each began with a 405 nm laser illuminating a non-linear crystal. Inside the crystal, a small fraction of the photons from the laser were converted into pairs of spatially-entangled lower-energy photons through the process of SPDC. The two photons in each pair produced through SPDC are generated concurrently, so they would be detected at the same time in detectors equidistant from the crystal. Each experiment includes either two or three detectors, labeled A, B, and B’. The photons arriving at these detectors were converted into electronic signals, which were sent through a set of electronics that generated the number of coincidence counts for all combinations of detectors. Coincidence counts are the number of times photons arrived at each of the detectors within the same very short time window, indicating that the detected photons were part of the same entangled pair. The single and coincidence counts were displayed on a LabVIEW computer interface. The lab guides provided to the students are detailed in Ref. [65] and summarized below.
@@
The second experiment was a continuation of the first. When the students arrived, the laser was already aligned through the crystal, and detectors A and B’ were already aligned with the pairs of entangled photons. The students began by placing a beam splitter in front of detector B’ to split the path of the photons going to that detector, and then added a fiber coupler for the third detector (detector B) to detect the photons that were reflected at the inserted beam splitter. The students then measured the two- and three-way coincidence counts and used those to calculate the second-order correlation parameter, , of the two outputs of the beam splitter. This allowed them to determine whether the light exiting the two outputs of the beam splitter was correlated (as it would be for a classical model of light, in which the amplitude of the wave splits) or anti-correlated (as it would be for single photons that can be detected at only one of the two outputs). The students obtained a value for the three-detector correlation parameter indicating that there were very few times all three detectors recorded counts at the same time, and thus that the experiment was best described by a quantum model of light. The students then made a measurement of the correlation parameter while ignoring detector A and obtained a value indicating a correlation between the outputs of the beam splitter, and thus a classical model of light. This was intended to demonstrate that there was a single-photon state only when the students accounted for the photons hitting detector A.
@@
In between the second and third experiments, V.B. added additional optical elements to create and align a Mach-Zehnder interferometer in one of the paths of the down-converted photons. The students began the third experiment by looking at interference with a visible laser (referred to as the “alignment laser”) that was already aligned along that same path. The students already had some familiarity with the alignment laser since they had also used it to help align the beam splitter in the second experiment. The students then spent the rest of the experiment working with the down-converted photons, with which they were immediately able to see interference. To better understand what they were seeing, the students had opportunities to play around with this setup by blocking and unblocking the two arms of the interferometer and ramping the piezoelectric actuator (piezo) that was attached to one of the mirrors in the interferometer. By using a computer-aided control system to apply different voltages to the piezo, the students were able to change the relative path length between the two arms of the interferometer by a distance on the nanometer scale, thereby causing the interference pattern to shift. To finish off the third experiment, the students placed half-wave plates in both arms of the interferometer and rotated one so that the polarizations of the photons in the two arms were orthogonal, thus removing the interference. When they then placed a polarizer (aligned at to the light in both arms of the interferometer) after the second beam splitter, the interference pattern reappeared. This is called a quantum eraser since the which-path information has been “erased.”
@@ III.2 Data sources
Our primary data source is think-aloud lab sessions where each pair of students worked together through a sequence of three two-hour lab sessions while being prompted to discuss their thinking out loud. During these sessions, V.B. acted as both a researcher (explaining the purpose and process of the research and prompting the students to explain their reasoning if needed) and an instructor or teaching assistant (answering questions the students had and ensuring they followed safety protocols). Since she had set-up the experimental apparatus and designed the lab guides, she was familiar with the experiments in a similar way to a typical instructor. V.B. tried to emphasize student reflection on the concepts covered in the lab and minimize prompting, so there were times when the students had long sense-making conversations on their own and other times when she asked the students questions to help them understand or explain what they were seeing, as an instructor might (see Appendix A for specific examples). All of the lab sessions were video and audio recorded.
After the students had completed all three experiments, they also participated in semi-structured individual interviews over Zoom, which we refer to as “post-interviews.” These lasted approximately one hour and included questions about the concepts covered in these experiments; where in the sequence of experiments the students had learned about these concepts; and other potential learning outcomes such as interest, self-efficacy, and feeling that they had seen real quantum effects [20]. We asked demographic questions at the end of the post-interviews and found that three of the students identified as white and the fourth as white and Asian 111The student specified a specific country in Asia when asked about their race, but to help preserve their anonymity, we have chosen to present only a more general racial categorization here. Two of the students identified as women and two as men, although we use gender neutral pseudonyms and pronouns throughout the paper to help protect their anonymity. The students were compensated with a gift card for the time they spent in the lab sessions and post-interview.
@@ III.3 Data analysis
Since the students spent the most time sense-making in the third experiment, we focused our analysis of resources solely on that experiment and therefore students’ understanding of single-photon interference. We began to identify student resources by watching the video recordings of the relevant conversations we had noted in the content logs. To avoid missing additional moments, we fully transcribed the third lab session for both groups. After carefully going through the transcripts and returning to watch the video clips of any moments where it seemed information could be missing (e.g., when students were using hand gestures), key conversations were chosen that demonstrated evidence of the identified resources. These were matched with student quotes from the post-interviews where the students reflected on the knowledge they had relied on while working with the experiments.
Our selection and naming of resources is based on the framework presented in Sec. II.2. We endeavored to choose names for the resources that were as close as possible to the students’ own words, avoiding any judgement about correctness, as the important part was finding student ideas that led to productive reasoning. Although resources can be any size, we are presenting here a set of “small-scale” resources that may be more in line with many PER studies since that is what instructors can most easily use [31, 40]. However, students often used many of these resources in conjunction with one another and they were not always easily distinguishable, so we also provide two larger categories of resources that we found useful when considering instructional implications.
@@ III.4 Limitations
Although we modeled our lab sessions after those used in actual courses, we investigated only one specific implementation of the single-photon experiments and it was in a non-class setting. The materials we designed were based off of materials used in other courses, but still had to be altered to fit within our specific constraints. The context of a research setting is different than a classroom since V.B. was with the students at all times (whereas teaching assistants or instructors often rotate through groups), the students did not receive additional instruction beyond the lab sessions (for example, detailed derivations of the math demonstrated in the experiment were only available if the students decided to look up the references in their own time), and the students did not need to consider grades or lab write-ups. Additionally, it is not possible to fully separate the impact of instruction from the impact of these experiments on their own, as the lab guides and interactions with the instructor are a part of students’ experiences with experiments. We therefore do not know exactly what affect V.B.’s interactions with the students had on their reasoning, although the interactions were intended to mimic that of an actual classroom environment as much as possible.
@@ IV Results: Student resources for understanding single-photon interference
To answer our first research question, we identify ideas students employ in the third lab session (the single-photon interferometer experiment). There, the students were compelled to make sense of the experimental data that simultaneously demonstrated particle-like and wave-like properties of single photons. This sense-making process consisted of the students discussing their observations with their lab partners and collectively building on various resources the students came in with until they had generated an explanation with which they were satisfied. Using the transcripts of the third think-aloud lab sessions, as well as the post-interviews, we identified resources the students utilized and list them in Table 1. We group some of the resources into broader categories because students often used several resources in the same category in similar ways in a given moment.
| Resource | Category |
| Waves can constructively and destructively interfere | Knowledge of classical wave interference |
| Need two things for interference to occur | Knowledge of classical wave interference |
| Things at same place and time can interfere | Knowledge of classical wave interference |
| Orthogonal things do not interfere | Knowledge of classical wave interference |
| Waves can split and recombine | Knowledge of classical wave interference |
| Knowledge of what laser interference looks like | Knowledge of classical wave interference |
| Wave functions represent probability distributions | Knowledge of wave functions |
| Wave functions can constructively and destructively interfere | Knowledge of wave functions |
| Wave functions have phases | Knowledge of wave functions |
| Wave functions have spatial components | Knowledge of wave functions |
| Wave functions have temporal components | Knowledge of wave functions |
| Quantum outcomes are probabilistic | - |
| Thinking in terms of information | - |